# HP-BERT: A framework for longitudinal study of Hinduphobia on social media via LLMs

### v1

@@ 6 Discussion
In ancient times, Hinduism was widespread and in the majority across many regions, but over centuries, political changes, invasions, and the spread of other religions led to a significant decline in the Hindu population in several countries [134, 135, 136]. In Pakistan [137, 138] Hinduism was once the dominant religion, especially in regions like Sindh and Punjab, with the Indus ValleyCivilisationn also having Hindu-related practices. However, where Hindus likely made up ($60\%-70\%$) of the population. Today, Hindus make up only ($1.85\%$) of Pakistan’s population [^10] https://en.wikipedia.org/wiki/Hinduism_in_Pakistan, which is due to forced conversions and low birth rates. In Bangladesh, Hinduism [139, 140] was once the dominant religion, especially under the Pala dynasty, with Hindus constituting around ($50\%-60\%$) of the population and now, Hindus make up about ($8\%-10\%$) of the population [^11] https://en.wikipedia.org/wiki/Hinduism_in_Bangladesh. In Afghanistan [141, 142], Hinduism was widely practised, especially in the Gandhara region, where Hindus were about ($40\%-50\%$) of the population, but today, only about ($0.01\%$) remain. Indonesia [143, 144] was also once predominantly Hindu, particularly during the Majapahit Empire, with Hindus making up around ($70\%-80\%$) of the population, but now only ($1.7\%$) of the population follows Hinduism, with Bali as the exception. The Maldives [145] was once a Hindu-majority region, with nearly ($90\%-100\%$) Hindus before the 12th century, but today, less than ($0.01\%$) of the population remains Hindu. In Sri Lanka [146, 147] Hinduism was widespread, especially in the Tamil-majority northern regions, where it was around ($30\%-40\%$) of the population in ancient times, but today, Hindus make up about 12% of the population. This shows how Hinduism, once dominant in these areas, has now become a minority religion in most of them.
@@
The lack of education about the scientific and philosophical heritage of Hinduism, both within India and globally, exacerbates the ridicule of its achievements. Some of the prominent Hindu philosophers such as Adi Shankaracharya (788–820 CE), Swami Vivekananda (1863–1902), Ramanuja (1017–1137 CE), and Madhvacharya (1238–1317 CE), along with mathematicians/scientists such as Aryabhata (476–550 CE), Srinivasa Ramanujan (1887–1920), Bhaskaracharya (Bhaskara II) (1114–1185 CE) made monumental contributions to philosophy, mathematics, astronomy, and science. Modern thinkers and educators, including Mahatma Gandhi (1869–1948) and Dr. B. R. Ambedkar (1891–1956) [150, 151] also significantly influenced intellectual and social thought. Despite their profound contributions, they have been underrepresented in educational curricula in India and abroad. This lack of recognition perpetuates a distorted narrative that dismisses ancient Indian knowledge, overshadowed by colonial biases. Internationally, the limited exposure to Hindu philosophy and science further reinforces stereotypes, creating a cultural vacuum filled with ridicule rather than respect.
This study has several limitations that should be acknowledged. First, all three datasets were primarily derived from X (Twitter), which may not fully capture the diversity of public opinion on Hinduphobia. The platform’s user base skews younger, more urban, and technologically savvy tweeters [152] with potential demographic biases. Additionally, the 280-character limit restricts the depth and context of expressed sentiments, making it difficult to capture nuanced perspectives. Wankhade et al. [153] reported that the informal language, abbreviations, and emojis further complicate sentiment analysis, reducing the precision of models such as our HP-BERT and HateBERT. These constraints highlight the need for incorporating additional data sources and methodologies to gain a broader understanding of Hinduphobia. Second, our HP-BERT model struggled with recognising sarcasm, irony, and negation, all critical for accurate sentiment interpretation. Nevertheless, sarcasm is also difficult to be recognised by humans and also has cultural biases that affect its understanding. It also faced challenges in identifying spam and fake content, which could undermine the reliability of classification. Despite efforts to filter and clean the dataset, noisy data such as irrelevant tweets or mislabelled sentiments remains an issue, distorting analysis and reducing accuracy.
@@
In future research, the scope of data sources can be expanded beyond Twitter to include other social media platforms such as Facebook, Instagram, Reddit, 4chan (known for extremist and controversial discussions), VK (a Russian platform with less moderation), and platforms such as Parler, Gab, and Weibo. Each of these platforms has distinct user demographics and moderation policies, making their inclusion valuable for capturing a broader spectrum of opinions and emotions. Incorporating data from these sources would provide a more comprehensive understanding of public sentiment, which can be extended to improving models for other problems such as Sinophobia [17]. Additionally, while BERT-based models have proven robust in sentiment analysis, future work could explore the application of more advanced large language models [155, 156]. As the internet is dominated by GenZ users (those born between the mid-1990s and early 2010s), who often use GIFs, emojis, and memes to express their opinions [157], we need to better account for non-textual and visual forms of communication. Humour, irony, and implicit messages in memes represent unique challenges in sentiment analysis that require innovative approaches. Future studies could utilise irony detection models and compare their performance with other sentiment analysis models. This comparison would help identify the strengths and limitations of various approaches, informing the development of more sophisticated tools. Longitudinal studies could also be conducted to analyze sentiments in entertainment media, such as movies and music, over the past few decades to identify trends and shifts in public sentiment. Finally, future sentiment analysis should integrate cultural, religious, historical, political, and economic factors, as these significantly influence public sentiment. For instance, during the COVID-19 pandemic, these factors played a pivotal role in shaping narratives and biases, including the rise of Hinduphobia. Events such as the portrayal of religious gatherings as superspreaders or the misrepresentation of traditional practices contributed to biased narratives. Therefore, future research can provide a more nuanced and comprehensive understanding of the drivers behind Hinduphobic sentiments, offering deeper insights into how such biases evolve and are sustained during global crises.
7 Conclusion
@@
The COVID-19 pandemic has significantly amplified Hinduphobic sentiments, resulting in widespread stereotyping and discrimination against the Hindu community. Our study employed LLMs to perform sentiment analysis on social media data, with a focus on Hinduphobic narratives during the pandemic. We curated a Hinduphobic COVID-19 X (Twitter) Dataset and fine-tuned the Hindupbobic-BERT model to review the evolution of Hinduphobic sentiments over time, including COVID-19. The HP-BERT model demonstrated moderate effectiveness with an additional sentiment analysis component using SenWave dataset.
@@
The HP-BERT model can be downloaded from Zenodo  [^13] https://zenodo.org/records/14607359.
@@
Table 9: Sentiment analysis of Hinduphobic discourse, showcasing positive polarity scores for a diverse sample of tweets from different countries.
@@
Table 10: Sentiment analysis of Hinduphobic discourse, showcasing negative polarity scores for a diverse sample of tweets from different countries.

### v2

@@ VI Discussion
In ancient times, Hinduism was widespread and in the majority across many regions, but over centuries, political changes, invasions, and the spread of other religions led to a significant decline in the Hindu population in several countries [142, 143, 144]. In Pakistan [145, 146], Hinduism was once the dominant religion, especially in regions like Sindh and Punjab, with the Indus Valley Civilisation also having Hindu-related practices. However, Hindus likely made up ($60\%-70\%$) of the population. Today, Hindus make up only ($1.85\%$) of Pakistan’s population [147], which is due to forced conversions and low birth rates. In Bangladesh, Hinduism [148, 149] was once the dominant religion, especially under the Pala dynasty, with Hindus constituting around ($50\%-60\%$) of the population, and now, Hindus make up about ($8\%-10\%$) of the population [150]. In Afghanistan [151, 152], Hinduism was widely practised, especially in the Gandhara region, where Hindus were about ($40\%-50\%$) of the population, but today, only about ($0.01\%$) remain. Indonesia [153, 154] was also once predominantly Hindu, particularly during the Majapahit Empire, with Hindus making up around ($70\%-80\%$) of the population, but now only ($1.7\%$) of the population follows Hinduism, with Bali as the exception. The Maldives [155] was once a Hindu-majority region, with nearly ($90\%-100\%$) Hindus before the 12th century, but today, less than ($0.01\%$) of the population remains Hindu. In Sri Lanka [156, 157] Hinduism was widespread, especially in the Tamil-majority northern regions, where it was around ($30\%-40\%$) of the population in ancient times, but today, Hindus make up about 12% of the population. This shows how Hinduism, once dominant in these areas, has now become a minority religion in most of them.
@@
The lack of education about the scientific and philosophical heritage of Hinduism, both within India and globally, exacerbates the ridicule of its achievements. Some of the prominent Hindu philosophers such as Adi Shankaracharya (788–820 CE), Swami Vivekananda (1863–1902), Ramanuja (1017–1137 CE), and Madhvacharya (1238–1317 CE), along with mathematicians/scientists such as Aryabhata (476–550 CE), Srinivasa Ramanujan (1887–1920), Bhaskaracharya (Bhaskara II) (1114–1185 CE) made monumental contributions to philosophy, mathematics, astronomy, and science. Modern thinkers and educators, including Mahatma Gandhi (1869–1948) and Dr. B. R. Ambedkar (1891–1956) [160, 161] also significantly influenced intellectual and social thought. Despite their profound contributions, they have been underrepresented in educational curricula in India and abroad. This lack of recognition perpetuates a distorted narrative that dismisses ancient Indian knowledge, overshadowed by colonial biases. Internationally, the limited exposure to Hindu philosophy and science further reinforces stereotypes, creating a cultural vacuum filled with ridicule rather than respect. These findings align with social identity threat theory, economic uncertainty, and health fears create conditions where out-group discrimination intensifies. However, our results extend beyond previous research on pandemic-driven xenophobia [17, 89] by documenting religion-specific patterns that persist across diverse national contexts. The moderate but consistent correlations across countries suggest universal mechanisms underlying crisis-driven religious discrimination.
This study has several limitations that should be acknowledged. First, all three datasets were primarily derived from X (Twitter), which may not fully capture the diversity of public opinion on Hinduphobia. The platform’s user base skews younger, more urban, and technologically savvy tweeters [162] with potential demographic biases. Additionally, the 280-character limit restricts the depth and context of expressed sentiments, making it difficult to capture nuanced perspectives. Wankhade et al. [163] reported that the informal language, abbreviations, and emojis further complicate sentiment analysis, reducing the precision of models such as our HP-BERT and HateBERT. These constraints highlight the need for incorporating additional data sources and methodologies to gain a broader understanding of Hinduphobia. Second, our HP-BERT model struggled with recognising sarcasm, irony, and negation, all critical for accurate sentiment interpretation. Nevertheless, sarcasm is also difficult to recognise by humans and has cultural biases that affect its understanding. It also faced challenges in identifying spam and fake content, which could undermine the reliability of classification. Despite efforts to filter and clean the dataset, noisy data, such as irrelevant tweets or mislabelled sentiments, remain an issue, distorting analysis and reducing accuracy.
@@
In future research, the scope of data sources can be expanded beyond Twitter to include other social media platforms such as Facebook, Instagram, Reddit, 4chan (known for extremist and controversial discussions), VK (a Russian platform with less moderation), and platforms such as Parler, Gab, and Weibo. Each of these platforms has distinct user demographics and moderation policies, making their inclusion valuable for capturing a broader spectrum of opinions and emotions. Incorporating data from these sources would provide a more comprehensive understanding of public sentiment, which can be extended to improving models for other problems, such as Sinophobia [17]. However, our current reliance on Twitter data highlights limitations that must be addressed to ensure generalizability.
We acknowledge several limitations in our study. First, while HP-BERT demonstrated strong performance across key metrics, its effectiveness is still constrained by the nature of the dataset, which may contain labeling inconsistencies and cultural nuances that are difficult to interpret. Additionally, although we provided a comparison with HateBERT, our evaluation primarily focused on a single labeled dataset (Hinduphobic COVID-19 X [165]), which may introduce biases due to its source (Twitter) and the time frame of data collection. These factors may limit generalizability to other platforms or events. We also recognize that visualised comparisons, while useful, do not fully capture deeper semantic nuances or potential classification overlaps. Moreover, although our results align with general theories in hate speech detection and previous findings on BERT-based model performance, alternative explanations, such as dataset skew or topic drift, could also contribute to the observed outcomes. Although our methodology involved both manual and GPT-assisted annotation to increase reliability, inter-annotator agreement was not formally measured, which could affect the consistency of ground truth labels. Future studies should address these concerns by conducting cross-platform validations, measuring annotator agreement, and expanding the datasets to include more diverse languages, dialects, and social contexts. This would allow a more robust evaluation of model generalizability and reduce the risk of overfitting to specific sociopolitical or temporal contexts. Future work could prioritize diverse datasets and robust evaluation methods to overcome these challenges. Additionally, while BERT-based models have proven robust in sentiment analysis and more advanced LLMs [166, 167].
@@
Although our framework focuses on BERT-based fine-tuning for Hinduphobic content detection, we acknowledge that advanced strategies could further enhance performance and flexibility. The techniques, such as block-based or modular neural architectures integrated with decision gates, present promising avenues for future exploration. In particular, CNN-based block designs [168] have shown effectiveness in tasks like anomaly detection, and similar ideas could be applied to fine-grained hate speech or sentiment analysis. Incorporating these modular and multi-block strategies may enable more robust feature extraction and task-specific adaptation. Building on these advancements, we must also account for non-textual forms of communication, as the internet is dominated by GenZ users (those born between the mid-1990s and early 2010s), who often use GIFs, emojis, and memes to express their opinions [169]. Humour, irony, and implicit messages in memes represent unique challenges in sentiment analysis that require innovative approaches. For instance, future studies could utilise irony detection models and compare their performance with other sentiment analysis models to identify strengths and limitations, informing the development of more sophisticated tools. Expanding beyond social media, longitudinal studies could analyze sentiments in entertainment media, such as movies and music, over decades to uncover trends in public sentiment, provided they address the data biases noted earlier. Finally, to deepen our understanding, sentiment analysis should integrate cultural, religious, historical, political, and economic factors, as these significantly shape public sentiment. The COVID-19 pandemic, for example, highlighted how such factors fueled Hinduphobia, with events such as the portrayal of religious gatherings as superspreaders or the misrepresentation of traditional practices that promote biased narratives. Therefore, future research can offer nuanced insights into how Hinduphobic sentiments evolve and persist during global crises.
VII Conclusion
@@
The COVID-19 pandemic has significantly amplified Hinduphobic sentiments, resulting in widespread stereotyping and discrimination against the Hindu community. Our study employed LLMs to perform sentiment analysis on social media data, with a focus on Hinduphobic narratives during the pandemic. We curated a Hinduphobic COVID-19 X (Twitter) Dataset and fine-tuned the Hinduphobic-BERT model to review the evolution of Hinduphobic sentiments over time, including COVID-19. The HP-BERT model demonstrated moderate effectiveness with an additional sentiment analysis component using the SenWave dataset.
@@
Hinduphobic COVID-19 X (Twitter) Dataset
The ”Hinduphobic COVID-19 X (Twitter) Dataset” can be accessed on Kaggle  [^8] https://www.kaggle.com/datasets/ashutoshsingh22102/hinduphobic-covid-19-x-twitter-dataset-india.
@@
The HP-BERT model can be downloaded from Zenodo  [^9] https://zenodo.org/records/14898138.
@@
TABLE XI: Sentiment analysis of Hinduphobic discourse, showcasing positive polarity scores for a diverse sample of tweets from different countries.
@@
TABLE XII: Sentiment analysis of Hinduphobic discourse, showcasing negative polarity scores for a diverse sample of tweets from different countries.
[…]